At the same rotational speed and coefficient of static friction between sock and the drum, would a sock of twice the mass slide down the wall? Explain your answer


1
Expert's answer
2021-11-21T17:31:05-0500

A wet socks clinging to the inside of a washing machine drum which is spinning at a speed of 4.7 m/s. The radius of the drum is 30 cm.



At the same rotational speed (V) and coefficient of static friction ("\mu_s" ) sock will not up down the wall if mass is increase to 2 m. Because "\mu_s" does not depend on mass it only depend on rotational speed as ("\mu_s = \frac{gR}{V^2}" )

If we write new equation for mass 2 m sock.

"N’ = \frac{2mV^2}{R} \\

\mu_s’ N’ ≥ 2mg \\

\mu_s’ \times \frac{2mV^2}{R} ≥ 2mg \\

\mu_s’ ≥ \frac{gR}{V^2} \\

\mu_s’ = \mu_s"

If rotational speed decrease then sock need more "\mu_s" to stick to wall.